OpenStudy (anonymous):

PLEASEE HELP. I NEED HELP FAST..PLEASE ;( Identify the 8th term of the geometric sequence -2, -12, -72 (2 points) -3,359,232 -559,872 559,872 3,359,232 Identify the 12th term of a geometric sequence where a1 = 8 and a6 = -8,192. (2 points) 134,217,728 33,554,432 -33,554,432 -134,217,728

OpenStudy (anonymous):

should be in math geometric sequence a1 = -2 a2 = -2(6) a3=-2(6)(6) = -2(6^2) an = (-2)6^(n-1)